1 – Assume that soon after graduation, you plan to start a new service providing business. (e.g. restaurant, lawn-care services, etc.) Propose an explicit service guarantee for your business. Clearly explain why you Included the features of your service guarantee.

2 – A certain gaming laptop manufacturer is re-designing the cooling system of a laptop model due to failures in cooling fans Currently, the cooling system uses one fan with the reliability of 0.95. The designers suggest to include two fans (in parallel) in the cooling system, instead of one, each with a reliability of 0.97. What percent improvement in reliability does adding the second fan provide?
Answer #1

1

The service guarantee may depend on the type of business you start. However, considering that every service has certain common criteria or factors on which they are often judged, we need to provide a service guarantee that is valuable. In this case, I would choose time as the service guarantee parameter. For example, a restaurant service can provide a service guarantee stating that all the dishes will be served in 15 minutes for in-house dining and 45 minutes for home delivery. For a lawn-care service we could provide a service guarantee that once the order is placed, we will reach the house to carry out the service within 4 hours.

2

Current reliability is 0.95. Now if we add two other fans in parallel each with a reliability of 0.97 then we need to first calculate their probability of failure (in case of parallel systems).

The total probability of failure is (1-0.95)*(1-0.97)*(1-0.97) = 0.000045.

This means the reliability of the system is 1-0.000045 = 0.999955

The percent improvement in reliability is (0.999955-0.95)/0.95 = 0.0525 or 5.25%

The improvement is 5.25%